MATH 136 – Calculus 2, section 1 Problem Set Information and Guidelines Spring 2014

Weekly homework assignments will be posted each Friday on the course homepage and collected the following Friday. Each assignment will consist of two parts

• an ‘A’ section to be completed on the WebAssign online homework system, and

• a ‘B’ section to be done on paper and handed in at the start of class. The ‘B’ selections will be human-graded with comments.

Here are some informational points to know and guidelines to follow while doing your assignments.

‘A’ section problems–WebAssign problems

• Almost all of the WebAssign problems will be keyed to problems in the text. You can tell the book version of the problem from the number displayed at the top of the problem in the assignment. For example, the problem marked SCalcCC4 1.7.004 is problem 4 from section 1.7 of Stewart’s Calculus (Concepts and Contexts version), 4th edition.

• You will get the most out of the WebAssign problems (and have a permanent record of what you did to consult later on) if you work out your answers legibly on paper first, check your work, and only then enter them into the WebAssign system.

• You will get three tries to submit correct work for each WebAssign problem. No further work can be submitted for credit after the second try.

• Each time you submit your work for grading, you can get immediate feedback about whether your answers are correct. In addition, after the assignment due date, you can see a correct solution or an explanation of how the correct answer is derived in WebAssign.

• If you believe that the most recent version of what you submitted to WebAssign is correct, but your answer was still marked wrong, you can send an email to Prof. Little with a request for “further review” of your answer. Points will be added to your problem set score to account for any correct solutions WebAssign could not recognize as correct.

• Many of the problems have randomized elements. This means that the questions you will get as part of your problem set will not be exactly the same as (but should still be equivalent in difficulty to) the problems other students in the class will do.

• Many problems have a “Practice Another Version” option that will let you enter a solution to a similar non-credit problem and view a correct solution. For your own benefit, do not abuse this feature. Don’t just blindly copy the steps of the correct solution without understanding what is going on!

‘B’ section – handwritten paper problems

• The ‘B’ section problems are intentionally more challenging, less computational problems that will require you to apply what you have learned in new situations and write up the mathematics in a careful way by hand. You may want to do all of the ‘A’ section problems in each book section first before trying the ‘B’ problems.

• Your work must be (easily) legible and organized. I reserve the right to return any excessively messy or disorganized work ungraded. You will get one chance, within one class meeting, to resubmit work for credit in a case like this. Write up the problems in order. Use only one side of the page. Don’t worry too much about saving paper: Please do not cram problems together in multiple columns on the page. Staple your assignment together.

• No credit will be given for just the answers to the ‘B’ problems. You must submit full solutions, showing all the steps necessary to justify your answer.

• Work on making your solutions clear and concise. Make good use of notation and diagrams.

• Leave space for the grader to write comments.

• Show enough detail that a confused student in the class could follow your solution. Don’t expect that the reader already knows how to solve the problem.

• When a problem asks for an explanation, you should provide one written in complete sentences. Proofread what you have said to be sure it makes sense.

• No “and then a miracle occurs” moments in solutions, please! If you are left with a gap in a calculation that you just can’t bridge, acknowledge it. The grader will indicate how to fill the gap, if appropriate, in comments.

• Start the assignment early (e.g. the day it is posted). This will allow you ample time to consult with me or the Calculus Workshop, or discuss things in a study group if you get stuck on some problems. If you start the assignment on the day before it is due, you will be very unhappy!

• If you work with other students to solve a problem, you must write up your final solutions independently. (Plan on taking scrap notes when you work out the problem first, and then writing it up in a clean form by yourself afterwards.) Add a note to your solution listing the other people you consulted.